OpenStudy (anonymous):

∫ −2 3 sin 2 (x)+cos 2 (x)dx

OpenStudy (anonymous):

∫ −2 3 sin 2 (x)+cos 2 (x)dx

OpenStudy (anonymous):

is this your integral? \[\int\limits-23(sin^2x+\cos^2x)dx\]or this?\[\int\limits-23\sin^2x+\cos^2xdx\]

OpenStudy (anonymous):

\[\int\limits_{-2}^{3}Sin ^{2}(x) + Cos ^{2}(x) dx\]

OpenStudy (anonymous):

[x]-2 to 3 3+2=5

OpenStudy (anonymous):

thanks again